### v1

# Numerical simulation of type II primordial black hole formation

Abstract
This study investigates the formation of primordial black holes (PBHs) resulting from extremely large amplitudes of initial fluctuations in a radiation-dominated universe. We find that, for a sufficiently large initial amplitude, the configuration of trapping horizons shows characteristic structure due to the existence of bifurcating trapping horizons. We call this structure of the trapping horizons “Type II PBH”, while the structure without a bifurcating trapping horizon “Type I PBH”, which is typically generated from a relatively small amplitude of the initial fluctuation. In Ref. [1], in the dust-dominated universe, the Type II PBH can be realized by the Type II initial fluctuation, which is characterized by a non-monotonic areal radius as a function of the radial coordinate (throat structure) in contrast with the standard case with a monotonic areal radius (Type I fluctuation). Our research reveals that a type II fluctuation does not necessarily result in a type II PBH in the case of the radiation fluid. We also find that for the initial amplitude well above the threshold value, the resulting PBH mass may either increase or decrease with the initial amplitude depending on its specific profile rather than its fluctuation type.
1 Introduction
Primordial black holes (PBHs) play a crucial role in exploring the early universe, particularly in investigating inhomogeneities whose scales are much smaller than those of the fluctuations probed by the cosmological microwave background. Since PBHs are remnants of non-linear inhomogeneity in the primordial universe following the inflationary era, they are unique probes of statistical properties of the small-scale primordial inhomogeneity. They are considered as one of the candidates for dark matter and can be detected as black hole binaries by gravitational wave interferometers. According to the standard scenario, PBHs were formed through the gravitational collapse of significantly high-density regions during the radiation-dominated era, which had been generated in the inflationary period [2, 3, 4].
The mechanism of PBH formation can be briefly understood by considering a model of the gravitational collapse in the flat Friedmann–Lemaître–Robertson–Walker (FLRW) universe (see, e.g., Refs. [5, 6]). A criterion for the collapse has been firstly proposed by Carr [7] as $\delta\gtrsim w=1/3$, where $\delta$ is the amplitude of the density perturbation, and $w$ is the coefficient of the assumed linear equation of state $p=w\rho$ with $p$ and $\rho$ being the pressure and the energy density, respectively. A more accurate threshold value has been explored by numerically analyzing general relativistic non-linear dynamics [8, 9, 10, 5, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
In cosmological perturbation theory, the growth of the perturbation is often described by the evolution of the curvature perturbation $\zeta$ which represents the growing adiabatic modes. To perform numerical simulations of PBH formation, one must establish the initial conditions for the spacetime geometry and the fluid in the system. This involves setting the initial profile of curvature fluctuations at the super-horizon scale before the horizon entry to be consistent with the assumption of the growing adiabatic modes. These analytical constructions are typically achieved using the gradient expansion techniques [9, 14, 21].
In a spherically symmetric system, we introduce a type II primordial fluctuation that is characterized by a non-monotonic behavior of the areal radius as a function of the radial coordinate $r$. The areal radius $R$ is often expressed as $R\propto re^{\zeta}$ with $\zeta$ and $r$ being the curvature perturbation and radial coordinate, respectively. Then, if $\partial_{r}\zeta<-1/r$, which can be realized by a sufficiently large amplitude of $\zeta$ with the scale fixed, the Type II fluctuation is generally recognized. These fluctuations exhibit a distinct neck-like structure, with a minimal areal radius at $\partial_{r}R=0$.
Kopp, Hofmann, and Weller (KHW) [1] investigated PBH formation associated with the type II fluctuations by employing the Lemaître–Tolman–Bondi (LTB) solution, which is the analytic solution for the dust fluid ($w=0$). They investigated the causal structure of the spacetime generated from a specific profile of the type II fluctuation. They explicitly showed that it describes a spacetime of BH formation, which they call ‘type II PBH.’ In addition, they revealed distinct properties of the type II PBH from ‘type I PBH,’ which arises from typical type I fluctuations with the amplitude over the threshold. We aim to investigate the dynamics and the spacetime structure generated from type II fluctuations in the radiation-dominated universe, namely, to clarify the impact of the existence of the radiation pressure.
When we consider the abundance of PBHs, the contribution from type II fluctuations is often neglected because they are highly unlikely to be generated statistically, and the abundance of PBHs originating from them is typically suppressed. However, as mentioned in Ref. [22], they can be significant when the statistics of primordial fluctuations are highly non-Gaussian. Moreover, in the context of a single-field inflationary model, it has been recently discovered in Ref. [23] that type II fluctuations surround the bubbles formed when the inflaton overshoots the barrier. As non-Gaussianities increase, the production of vacuum bubbles becomes dominant, leading to a mass function and abundance primarily driven by these fluctuations. In addition, a similar black hole formation mechanism due to the birth of a baby universe with a tunneling process during inflation has been reported in Refs. [24, 25]. The baby universe is connected to our universe through a throat inside the black hole, like the neck structure in the type II fluctuation.
In the Misner–Sharp formalism [26] (see also Refs. [1, 27]), the evolution equations suffer from the trouble associated with $/0$ 0 term at the neck radius satisfying $\partial_{r}R=0$ when we consider type II fluctuations. To overcome this problem, it is necessary to conveniently use and modify the equations, as done in Ref. [24] for another scenario of PBH formation from the collapse of vacuum bubbles. On the other hand, in this paper, we use the simulation code developed from a 3+1 dimensional simulation code [28] with BSSN [29, 30] formalism with CARTOON method [31], so that we can avoid the same problem.
This paper is organized as follows. Section 2 describes the setup and numerical code of spherically symmetric numerical simulation of PBH formation. In Section 3.1, we present the simulation results, highlighting the time evolution of fluctuations. Then, we explore the formation of trapping horizons, including the BH horizon in Section 3.2, and give an interpretation of the resultant trapping horizon configurations in Section 3.3. In Section 4, we examine the evolution of PBH mass. In Section 4.2, to check the profile dependence of the PBH mass, we explore PBH formation from another initial profile that models the overdense region, as outlined by KHW. Finally, we offer our conclusions in the last section.
Throughout the paper, we use the geometrical unit in which the speed of light and the gravitational Newton constant are unity, i.e., $c=G=1$.
2 Initial data setting
We start our simulation from the time when the length scale of the fluctuation is much larger than the Hubble length $H_{\text{b}}^{-1}$, that is, $k^{-1}\gg(aH_{\text{b}})^{-1}$ with $k$, $a=a(t)$ and $H_{\text{b}}:=(da/dt)/a$ being the characteristic comoving wave number, the scale factor and the Hubble parameter of the background flat FLRW universe, respectively. We obtain the initial condition by applying the cosmological long-wavelength approximation, in which the small parameter $\epsilon:=k/(aH_{\rm b})$ is introduced, and the leading order spatial metric is given by [32]
$\gamma_{ij}=a^{2}(t)e^{2\zeta(x^{k})}f_{ij},\,$ (2.1)
with $f_{ij}$ being the reference 3-metric of the background flat geometry, and $\partial_{i}\zeta/(aH_{\rm b})=\mathcal{O}(\epsilon)$. We note that the function $\zeta$ is a time-independent arbitrary function of the spatial coordinates $x^{k}$. Once we initially set $\zeta$ as a function of the spatial coordinates $x^{i}$ and fix the gauge conditions, we can successively obtain the growing mode solution up to the following leading terms in this expansion [9, 21]. These long-wavelength solutions are used as the initial data for the time evolution. In this paper, we initially fix the gauge by imposing the constant mean curvature slice (uniform Hubble slice) and the normal threading, for which the trace of the extrinsic curvature of the time slice is uniformly given by $-3H_{\rm b}$ and the shift vector is equal to zero.
In a spherically symmetric system, the line element at the lowest order is denoted as
$ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (2.2)
where $d\Omega^{2}$ represents the line element of a 2-sphere. The areal radius is given by $R=are^{\zeta}$. Throughout this paper, we use the scale-up coordinate $z$, which is related to $r$ as
$r=z-\frac{\eta}{1+\eta}\frac{L}{\pi}\sin\left(\frac{\pi}{L}z\right)$ (2.3)
with a parameter $\eta=20$ (see Ref. [28] for details). In this coordinate, the resolution in the central region is $1+2\eta$ times finer than that around the boundary $z=r=L$.
Up through section 4, we use the following Gaussian-shape profile in the numerical simulation,
$\zeta(r)=\mu e^{-\frac{1}{2}k^{2}r^{2}}W(r),$ (2.4)
where $\mu$ is the amplitude of the fluctuation and $W(r)$ is a window function introduced to eliminate the tail of the Gaussian function and impose $\zeta=0$ at the outer boundary of the numerical domain (see Ref. [28] for the explicit functional form). Figure 1 shows the functional form of this profile (left) and the areal radius $R$ (right). By taking a sufficiently large amplitude of the fluctuation $\mu$, the areal radius $R(r)$ of the perturbed region becomes non-monotonic along the radial coordinate $r$, i.e., a pair of the local minimum and maximum points appears with $\partial_{r}R=0$ for $\mu\gtrsim 1.4$ as in Fig. 1. We call this structure neck or throat structure, and KHW defined such fluctuations as of type II and distinguished them from type I fluctuations, which have a monotonic areal radius [1].
Figure 1: Initial profile of the curvature fluctuation $\zeta$ (left) and areal radius $R$ as functions of the scale-up coordinate radius $z/L$. By taking the amplitude parameter $\mu$ larger, a type II fluctuation appears for $\mu=1.4$ and above with the neck $\partial_{r}R=0$.
The feature of type II fluctuations can also be understood as two peaks [23] of the Shibata–Sasaki compaction function $\mathcal{C}_{\text{SS}}$ defined by
$\mathcal{C}_{\text{SS}}(r)=\frac{1}{2}\left[1-\quantity(1+r\frac{d\zeta}{dr})^{2}\right]=\frac{1}{2}\left[1-\left(a^{-1}e^{-\zeta}\partial_{r}R\right)^{2}\right].$ (2.5)
This function has been firstly introduced in Ref. [9] as an estimator for the criterion of PBH formation, and its usefulness has been recognized through subsequent related works (see Ref. [33] and references therein for details including its relations to legitimate compaction functions in different gauges and physical interpretations). The outer peak of $\mathcal{C}_{\text{SS}}$ corresponds to the neck in the type II fluctuation, the radius at which the areal radius takes the minimum value considered relevant for the PBH formation. We may also consider a two-peak configuration with type I fluctuations [34] while we only focus on two peak configurations associated with type II fluctuations. The form of the compaction function for the specific function of $\zeta$ (2.4) is shown in Fig. 2. By using position of the peak of $\mathcal{C}_{\text{SS}}(r)$, $r_{\text{m}}=\sqrt{2}/k$, we analytically obtain the threathold value $\mu=e/2=1.359\cdots$.
Figure 2: Compaction function $\mathcal{C}_{\text{SS}}$. The horizontal axes are the scale-up coordinate radius $z/L$ (left) and the areal radius $R$ (right). By taking the amplitude $\mu$ larger, the number of peaks becomes two for $\mu\simeq 1.4$ and above, corresponding to type II fluctuations.
To follow the time evolution of gravitational collapse from the initial profile of fluctuation, we use the numerical code COSMOS-S [35, 36, 37] based on a 3+1 dimensional code of fourth order Runge–Kutta method with BSSN formalism [29, 30]. This code is specialized for spherically symmetric systems using the CARTOON method [31]. We consider the time evolution of the initial fluctuation given by the functional form of $\zeta$ (2.4) with the equation of state parameter $w=1/3$, namely, radiation fluid. Here, the asymptotically spatially flat FLRW boundary condition is implemented as in Ref. [9]. However, we could not avoid the constraint violation at the outer boundary, and this propagates into the bulk region. Therefore, in the following analyses, we terminate the time evolution before the violation reaches the central region relevant for the analyses of PBH formation.
We also find the constraint violation near the center for cases where an apparent horizon forms. Nevertheless, since we find both independent future-directed null directions are inward there, the constraint violation does not propagate outward. In practice, to avoid a crash of numerical calculation, we excise several grids inside the apparent horizon if needed.
We take the coordinate size $L$ of the numerical region ($0\leq z\leq L$ or, equivalently, $0\leq r\leq L$) as the unit length scale and set the initial value of the scale factor to unity. We fix the ratio between the background initial Hubble length $H_{\rm b}^{-1}$ and the typical scale $k^{-1}$ of the initial fluctuation as $k^{-1}=10H_{\text{b}}^{-1}$ which ensure the long-wavelength approximation. For a longer duration of the time evolution, we need to take a larger region than the initial Hubble length to avoid the propagation of the constraint violation. In the following, we take $H_{\text{b}}^{-1}=0.01L$ or $H_{\text{b}}^{-1}=0.005L$ depending on the duration of the required time evolution. The horizon entry time $t_{H}$ is defined as $k^{-1}=(aH_{\text{b}})^{-1}|_{t=t_{H}}$ throughout this paper.
3 Spacetime structure and a new definition of type I/II PBHs
3.1 Time evolution of type I/II fluctuations
First, we show the time evolutions of the fluid energy density and the lapse function in Fig. 3 for $\mu=0.5$, $1.2$, and $1.8$. For the cases of type I fluctuations $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig. 3. The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig. 3 as Ref. [37]. For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
(a) $\mu=0.5$
(b) $\mu=1.2$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ of type I fluctuations (left and middle) and a type II fluctuation (right). The initial configurations ($t=0.01t_{H}$) are shown in black on a different scale.
To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint (Fig. 4). One can find roughly quadratic convergence in the different spatial resolutions, and it is consistent with our numerical code with a second-order scheme implemented in evaluating the fluid flux.
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for three typical cases for no collapse (left), collapse from a type I fluctuation (middle), and collapse from a type II fluctuation (right), respectively. The horizontal axis is a time coordinate normalized by the time of the horizon entry. We use the simulation result to $t=50t_{H}$ to evaluate spacetime structures around horizon formation before the violation propagation reaches the bulk region from the outer boundary. For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH horizon (or the future outer trapping horizon, described in Sec. 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
3.2 Identification of BH horizon
First, let us introduce two independent future-directed radial null vector fields $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [26, 38, 39, 6].
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted. For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one. On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time. The worldlines of the intersections are shown in Fig. 6. In order to identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future-trapped and past-trapped regions, respectively. More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, we can identify the outermost boundary surface of the future-trapped region as an apparent horizon. The cosmological horizon would be the outermost boundary between the past-trapped and non-trapped regions. Still, the following section will present a more concrete classification of trapping horizons.
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$. For the $\mu=1.2$ case, the trapping horizon at the center is formed at around $t=40t_{H}$ (middle). For the $\mu=1.8$ case, the horizon is formed at around $t=24t_{H}$ (right).
(a) $\mu=0.5$
(b) $\mu=1.2$
(a) $\mu=0.5$
(b) $\mu=1.2$
(c) $\mu=1.8$
Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$. It corresponds to the points $(t,r)$ of the compactness $12M/R=$ 1 in Fig. 5.
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves. In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=$ 8.1 to 8.9.
3.3 Spacetime structures of type I/II PBH
Let us introduce the classification of trapping horizons following Ref. [39, 40]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when we consider a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of trapping horizons can be summarized as Table 1.
$\theta_{+}=0$ + 0 -
$\theta_{-}$ Past Bifurcating Future
$\mathcal{L}_{-}\theta_{+}$ Inner Degenerate Outer
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [39]. A black hole, white hole, and cosmological horizons with $w=1/3$ $(w=0)$ are identified with a future outer, past outer, and past degenerate (past inner) trapping horizons, respectively.
(a) Type I PBH
(a) Type I PBH
(b) Type II PBH
Figure 9: Sketch of type I and II PBHs in radiation domination. The left figure defines a type I PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type II PBH.
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs. 7 and 8 for the radiation case. From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 9 as possible conformal diagrams of type I and type II PBHs, respectively. The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively. See Figs. 16, 17 and 19 and Ref. [40]. Then, we confirm that the outermost boundary surface of the future-trapped region is the future outer trapping horizon, which describes an apparent horizon. In addition, one can find that the outermost boundary surface between the past-trapped and non-trapped regions describes the cosmological horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This crucial difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II nevertheless. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
It is worth noting that the schematic diagram given by Fig. 9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.
Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.
4 PBH mass
4.1 PBH mass with the Gaussian initial fluctuation
Let us consider how the mass of the PBH depends on the initial amplitude $\mu$ of the fluctuation in the type II regime by using the three-zone model with dust as seen in Fig. 18. The PBH mass is then given by the mass parameter for the Schwarzschild spacetime in the middle patch. If we fix the scale of the spatial curvature of the closed FLRW region, the horizon area of the Schwarzschild region is fixed by the horizon area on the matching surface between the closed FLRW and the Schwarzschild regions. Increasing the initial amplitude, we expect that the closed FLRW region will be extended and the boundary will be shifted more to the right. In the closed FLRW side, the shift of the boundary to the right results in the decreasing area of the horizons. Therefore, the mass parameter for the Schwarzschild region is also expected to decrease. From the above consideration of the model with dust, we may naively guess that the mass of the resultant PBH would decrease with the initial amplitude value for the type II regime. See also Ref. [41]. However, as will be shown soon, this guess is not necessarily true for the radiation case, and the behavior also strongly depends on the profile of the initial fluctuation.
Here, we evaluate the mass of the PBH in two different ways at each characteristic time. The one is to calculate the Misner-Sharp mass inside the apparent horizon at its first appearance with the future-trapped region (see Fig. 6). We refer to the mass at this time as the initial black hole mass. The apparent horizon appears as a future outer trapping horizon in a type I PBH, while it appears as a bifurcating outer trapping horizon in type II. We also calculate the final PBH mass by estimating the effect of the accretion. It is known that the accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [2, 42]. In this accretion model, the time evolution of the mass is given by
[…]
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion. But it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
$M(t)=\frac{1}{\frac{1}{M_{f}}+\frac{3F}{2t}},$ (4.2)
where $M_{f}$ is the final BH mass. We fix the values of $M_{f}$ and $F$ by fitting Eq. (4.2) to the numerical data in a sufficiently late finite time interval denoted in Fig. 10 together with the time evolution of the mass for each value of $\mu$. The values of $F$ obtained by the fitting are listed in Table 2. The value of $F$ decreases with increasing $\mu$, which means the accretion is getting smaller. This behavior is consistent with the results in Ref. [43].
Figure 10: The time evolution of the BH mass for each value of $\mu$. The black dashed curves are fitting curves with Eq. (4.2) including the two parameters $M_{f}$ and $F$. The fitting region is taken for the interval between $t=450t_{H}$ and $t=500t_{H}$.
$\mu$ $M_{f}[M_{\text{H}}]$ $F$
1.2 11.9 5.0
1.4 14.7 5.0
1.6 16.8 4.9
1.8 18.6 4.7
2.0 20.1 4.5
2.2 21.4 4.4
Table 2: The values of $F$ and $M_{f}$ for each amplitude of $\mu$ result from the fitting in Fig. 10.
Let us check the $\mu$ dependence of the PBH mass. In Fig. 11, we plot the initial and final PBH mass values for each value of $\mu$.
Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$. In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs. 6 and 9.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
4.2 PBH mass with the initial fluctuation motivated by the three-zone model
To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
$\zeta_{\text{cFLRW}}(r)=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}\quad\text{for}\quad r\leq r_{a}:=\frac{1}{k}\sin(k\chi_{a}),$ (4.3)
where $1/k$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
$\zeta(r)$ $=2\tilde{\mu}\ln\quantity(\Psi_{3}(r)-\Psi_{3}(L))W(r),$ (4.4)
$\Psi_{3}(r)$ $\coloneqq e^{\zeta_{\text{cFLRW}}(r)/2}+e^{\zeta_{\text{ref}}(r)/2},$ (4.5)
where $\zeta_{\text{ref}}\coloneqq 0.001e^{-\frac{1}{2}k^{2}r^{2}}$ and $\tilde{\mu}\coloneqq k\chi_{a}/\pi$. The overall factor $\tilde{\mu}$ is multiplied for convenience to realize the background flat FLRW universe in the limit $\tilde{\mu}=0$. The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$.
Figure 12: Functional form of $\zeta_{\text{cFLRW}}$ for overdense region $\chi\leq\chi_{a}$. This profile is obtained by imposing a boundary condition that $\zeta=0$ at the edge of the overdense region $\chi=\chi_{a}$ [1].
The initial profiles of $\zeta_{\rm cFLRW}$, $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 12, 13 and 14. This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 13: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$. By taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$.
Figure 14: Compaction function $\mathcal{C}_{\text{SS}}(r)$ for thnon-linearprofile. The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right).By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 15, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of nonlinear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 15.
Figure 15: Initial PBH mass with respect to the amplitude of fluctuation for the three-zone profile.
The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius. The initial PBH mass would be roughly estimated by the areal radius of the peak of the compaction function at the horizon entry. Therefore, if the areal radius around the peak of the compaction function increases and decreases with the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease, respectively. In the case of the Gaussian profile, one can find that the areal radius around the peak of the compaction function increases with the value of the amplitude $\mu$ as the right panel of Fig. 2, while it decreases with $\tilde{\mu}$ in the case of the specific profile (4.4) as the right one of Fig. 14.
5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH and a type II fluctuation does in a type II PBH. However, in the radiation fluid case, this is not true. We found a gap between the classifications in terms of the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs in terms of the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 43]. In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
Acknowledgments
The authors are grateful to K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). T.H. is supported in part by JSPS KAKENHI Grant Nos. 19K03876 and 20H05853. D.S. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. C.Y. is supported in part by JSPS KAKENHI Grant Nos. 20H05850 and 20H05853. C.Y. and T.H. acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [5], the spacetime of PBH is described by the patchwork of the three different regions : the innermost overdense region given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with nonlinear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. In spite of this great simplification, the model provides us with a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 16, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;*)(+;*)( + ; * ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH horizon and cosmological horizon become future outer (;)(-;-)( - ; - ) and past inner (+;+)(+;+)( + ; + ), denoted by red null and blue timelike lines, respectively.
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 16, and the resultant spacetime diagram is shown in the left panel of Fig. 18. In these conformal diagrams, we define the right and left directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
For the type II fluctuation, the extracted regions are yellow-shaded in Fig. 17. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for the type II is given by the right panel of Fig. 18.
(a) closed FLRW for dust
(b) Schwarzschild
(a) closed FLRW for dust
(b) Schwarzschild
(c) flat FLRW for dust
Figure 16: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 18 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past outer (+;)(+;-)( + ; - ), future inner (;+)(-;+)( - ; + ), and past inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red null, blue null, red timelike, and blue timelike lines, respectively.
(a) closed FLRW for dust
(b) Schwarzschild
(a) closed FLRW for dust
(b) Schwarzschild
(c) flat FLRW for dust
Figure 17: Same as Fig. 16 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
(a) Type I PBH
(a) Type I PBH
(b) Type II PBH
Figure 18: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I PBH as a causal structure of typical case PBH. The right one is constructed by changing the cutting diagram range and defining a type II PBH.
(a) closed FLRW for w=1/3
(a) closed FLRW for w=1/3
(b) flat FLRW for w=1/3
Figure 19: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null.
Appendix B Derivation of $\zeta_{\text{cFLRW}}$ in radial coordinate $r$
The leading order of the long-wavelength solution in terms of spatial 3-metric is expressed as
$dl^{2}=a^{2}e^{2\zeta(r)}\quantity(dr^{2}+r^{2}d\Omega^{2}),$ (B.1)
where $r$ is a radial coordinate in the conformally flat geometry. We consider this perturbed metric of an overdense region as closed FLRW geometry with a constant positive curvature $k^{2}$,
$dl^{2}$ $=a^{2}\quantity(d\chi^{2}+\frac{1}{k^{2}}\sin^{2}k\chi\,d\Omega^{2}),$ (B.2)
where $\chi$ means the comoving radial coordinate. By comparing the correspondence of the metric, we obtain
$d\chi$ =±$=\pm dre^{\zeta},$ (B.3)
$\frac{1}{k}\sin k\chi$ $=re^{\zeta},$ (B.4)
then we get the relation,
$R$ $=\frac{a}{k}\sin k\chi=are^{\zeta},$ (B.5)
$r$ $=C\exp\int\frac{d(k\chi)}{\sin k\chi}=C\tan\frac{k\chi}{2},$ (B.6)
where C is an integral constant. By using these relations, $\zeta$ is obtained as
$e^{\zeta}$ $=\frac{\sin k\chi}{kr},$ (B.7)
$=\frac{2}{Ck}\cos^{2}\frac{k\chi}{2},$ (B.8)
$=\frac{2}{k}\frac{1/C}{1+\quantity(r/C)^{2}}.$ (B.9)
If we write $C$ as $2C^{\prime}/k$ for simplicity,
$e^{\zeta}=\frac{1/C^{\prime}}{1+\frac{k^{2}}{4}\quantity(r/C^{\prime})^{2}}\quad\therefore\quad\zeta=\ln\frac{1/C^{\prime}}{1+\frac{k^{2}}{4}\quantity(r/C^{\prime})^{2}}.$ (B.10)
Here we impose the boundary condition that $\zeta=0$ for $\chi=\chi_{a}$, we get
$C^{\prime}=\cos^{2}\frac{k\chi_{a}}{2}.$ (B.11)
Then
$\zeta=2\ln\frac{\cos k\chi_{a}/2}{\sqrt{\cos^{4}k\chi_{a}/2+\frac{1}{4}k^{2}r^{2}}}.$ (B.12)